Match the following aqueous solutions with the appropriate letter from the column on the right. Assume complete dissociation of electrolytes.

1. 0.21 m NH4NO3 A. Lowest freezing point
2. 0.12 m CrCl2 B. Second lowest freezing point
3. 0.17 m NH4Br C. Third lowest freezing point
4. 0.39 m Sucrose (nonelectrolyte) D. Highest freezing point
Answer #1

Lower the value of ∆T will have lower freezing point.

∆T = i*Kf* molarity

Kf ( water ) = 1.86 °C/m

1. NH4NO3 , i=2 ( NH4+ & NO3-)

∆T = 2 * 1.86 * 0.21

= 0.7812

2. CrCl2 , i=3 ( Cr2+ & 2 Cl-)

∆T = 3*1.86 * 0.12

= 0.6696

3. NH4Br , i = 2 ( NH4+ & Br-)

∆T = 2* 1.86 * 0.17

= 0.6324

4. Sucrose , i=1

∆T = 1*1.86 * 0.39

= 0.7254

So in increasing order

NH4NO3 , sucrose , CrCl2 , NH4Br

If you have any questions please comment

If you satisfied with the solution please rate it thanks